# Does it matter? A more careful treatment of density fluctuations in 21-cm simulations

### v1

@@ Acknowledgements.
It is our pleasure to thank Julian B. Muñoz for reviewing the manuscript and sharing his comments. We would also like to thank Andrei Mesinger and Tal Adi for useful discussions. We also acknowledge the efforts of the 21cmFAST and CLASS authors to produce state-of-the-art public 21-cm and CMB codes. JF is supported by the Zin fellowship awarded by the BGU Kreitmann School. SL is supported by an Azrieli International Postdoctoral Fellowship. EDK acknowledges joint support from the U.S.-Israel Bi-national Science Foundation (BSF, grant No. 2022743) and the U.S. National Science Foundation (NSF, grant No. 2307354), as well as support from the ISF-NSFC joint research program (grant No. 3156/23).
Appendix A Scale-dependent matter evolution
Figure 11: Top panels: The ratio of $\sigma\left(M_{h},z\right)$ and the Sheth-Tormen HMF, when they are calculated with the SIGF and the SDGF. Bottom left panel: The ratio of the collapsed fraction of halos that host ACGs (with $M_{\mathrm{min}}$ given by Eq. (IV)), when calculated with the SIGF and the SDGF. Bottom right panel: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the matter over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
As we discussed in Sec. III.2, the local SFRD can be computed with the EPS formalism, where the local linear matter over-density $\delta_{m}^{\mathrm{lin}}\left(\mathbf{x},z\right)$ is considered. This quantity can be computed in Fourier space via $\delta^{\mathrm{lin}}_{m}\left(\mathbf{k},z\right)\approx\mathcal{D}_{m}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where
$\sigma^{2}\left(M_{h},z\right)=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\mathcal{T}^{2}_{m}\left(k,z=0\right)\mathcal{D}_{m}^{2}\left(k,z\right)W_{\mathrm{ball}}^{2}\left(kR_{M_{h}}\right).$ (26)
Appendix B Boltzmann equation for the spin temperature
@@
Combining Eqs. (B)-(37) thus yields
@@
Appendix C Temperature fluctuations in the dark ages
where $\kappa_{1-0}^{\mathrm{HH}}$ is the collision rate (in units of $\mathrm{cm^{3}/sec}$) of hydrogen atoms with themselves. Here, $\delta_{T}\left(k,z\right)$ and $\delta_{21,\mathrm{iso}}\left(k,z\right)$ are the Fourier transform of the fractional fluctuation in $T_{k}\left(\mathbf{x},z\right)$ and $T_{21}\left(\mathbf{x},z\right)$, respectively, and the latter does not contain fluctuations sourced from peculiar velocity (see Eq. (12)). At the dark ages, the fluctuations $\delta_{T}$ are sourced by the perturbations in the baryon density field $\delta_{b}$, and thus it is useful to define the so called “adiabatic index”,
$k_{0}-\Delta k/2<|\mathbf{k}|<k_{0}+\Delta k/2,$ (45)
Appendix D Scale-dependent Lagrangian perturbation theory
@@
D.1 First order - Zel’dovich approximation
@@
where we used the linear expression for $\delta_{i}$ (Eq. (9)) as we seek first order solution to $S_{i}$. To see that this solution is consistent, we expand the Euler-Poisson equation, Eq. (53), to first order,
@@
which is indeed the equation of motion for $\delta_{i}$ in linear perturbation theory, above the Jeans scale (c.f. Eq. (7)). Furthermore, the homogeneity assumption we made in Eq. (52) implies that $\delta_{i}$ and $S_{i}^{(1)}$ have the same zero initial conditions at $z=z_{0}$. Thus, the solution to $\mathbf{s}_{i}^{\left(1\right)}$ in Fourier space reads
@@ D.2 Second order - 2LPT
If variable separation was possible, then from Eq. (D.2) we could have obtained two different equations, one for the spatial part of $S_{i}^{(2)}$, and one for its temporal part. However, as we established in Eq. (58), variable separation cannot be performed for $S_{i}^{(1)}$, and thus $S_{i}^{(2)}$ cannot be separated to its spatial and temporal components either.
Note that given $S_{i}^{(1)}$, Eq. (D.2) does not give a differential equation for $S_{i}^{(2)}$, but rather for the Laplacian of $S_{i}^{(2)}$. In order to obtain an integro-differential equation for $S_{i}^{(2)}$, we take the Fourier transform of Eq. (D.2). While applying Fourier transform on the LHS gives a multiplication by $-k^{2}$, taking the Fourier transform of the RHS, being proportional to the square of second derivatives of $S_{i}^{(1)}$, would result in a convolution. We use the fact that in Fourier space $\ddot{S}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)+2H\dot{S}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)=4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(1\right)}\left(\mathbf{k},z\right)$ and apply the inverse Fourier transform to arrive at
@@
D.2.1 Scale independent growth
While in principle Eq. (D.2) can be solved numerically, this is computationally expensive, as for each cell $\mathbf{q}$ in the grid, double integration (that involves all the cells in the box in Fourier space) is required, for both species. To ease the computational cost, we will make the approximation that is done vastly in the literature, of approximating $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$, i.e., we replace the SDGF for each species with a universal SIGF. Then, from Eq. (58), we write
@@ D.2.1 Scale independent growth
where $\psi^{\left(1\right)}\left(\mathbf{q}\right)$ is the inverse Fourier transform of $\delta_{0}\left(\mathbf{k}\right)/k^{2}$ (or equivalently, in real space $\delta_{0}=-\nabla^{2}_{\mathbf{q}}\psi^{(1)}$). Because $\psi^{(1)}$ is derived from $\delta_{0}$, which is common for all species, under the SIGF approximation the solution to $S_{i}^{\left(1\right)}$ is a universal (i.e. independent of the species $i$) solution $S^{\left(1\right)}$. Plugging this universal solution back to Eq. (D.2) yields
@@
Because the RHS of Eq. (D.2.1) is universal for all species, and since $S_{i}^{(2)}$ has zero initial conditions at $z=z_{0}$ for every species $i$, we deduce that the SIGF approximation implies that $S_{i}^{(2)}\left(\mathbf{q},z\right)$ also has a universal solution $S^{(2)}\left(\mathbf{q},z\right)$. This property can be further inferred for any higher order in perturbation theory. Hence, the SIGF approximation corresponds to solving the perturbed equations of motions under the approximation $\delta_{b}\approx\delta_{c}\approx\delta_{m}$, namely we directly solve for the total matter density field.
After substituting the universal solution $S_{i}^{(2)}\left(\mathbf{q},z\right)=S^{(2)}\left(\mathbf{q},z\right)$, Eq. (D.2.1) can be written in the following form,
@@
D.2.2 Solution to $E\left(z\right)$
@@
Figure 14: Left panel: realization of the baryon over-density field at $z=20$, evolved non-linearly with scale-dependent 2LPT. Right panel: the fractional error in the normalized density field ($\Delta_{i}\equiv 1+\delta_{i}$) due to approximating $\delta_{b}\approx\delta_{c}$.
Figure 15: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$ (solid curves), $k=0.3\,\mathrm{Mpc}^{-1}$ (dashed curves), and $k=0.15\,\mathrm{Mpc}^{-1}$ (dotted curves), when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). For fair comparison, in all curves we used the non-linear $\delta_{m}^{\mathrm{NL}}$ in the EPS formalism, as in standard 21cmFAST.
D.3 The effect of scale-dependent growth on the 21-cm signal at cosmic dawn
In 21cmFAST, the SIGF approximation of $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$ is taken, and Eqs. (62), (74) are implemented in the code for computing the first and second order expressions for the displacement vector field. In order to compute the non-linear baryon density field more precisely, we replaced Eq. (62) in 21cmFirstCLASS with Eq. (61), i.e. we compute the first order expression for $\mathbf{s}_{i}^{(1)}$ via the SDGF $\mathcal{D}_{i}\left(\mathbf{k},z\right)$, for each species $i$. As for the second order expression, $\mathbf{s}_{i}^{(2)}$, we did not attempt to solve the coupled integro-differential equations given by Eq. (D.2), and instead we keep using the approximated Eq. (74). This is justified as by construction the second order correction $\mathbf{s}_{i}^{(2)}$ has a smaller contribution to $\mathbf{s}_{i}$, compared to the first order correction $\mathbf{s}_{i}^{(1)}$. We leave for future work the comparison of the approximated analytical solution of Eq. (74) with the fully numerical solution of Eq. (D.2).
We show in Fig. 14 the comparison between the more precise $\delta_{b}^{\mathrm{NL}}$, computed with the scale-dependent first order LPT, and the common calculation that evolves the density field with the SIGF. At $z=20$, we see that the error is of the order of a percent at the most over- and under-dense regions. At higher redshifts, this error is expected to increase at higher redshifts due to the departure of the SDGF from the SIGF, while at lower redshifts it vanishes as $\mathcal{D}_{b}\left(k,z\right)\to D\left(z\right)$, see Fig. 3.
The error in the evaluation of $\delta_{b}^{\mathrm{NL}}$ with the SIGF is enhanced when the 21-cm power spectrum is considered, due to the intricate interplay between non-linear fluctuations in different fields, e.g. $T_{k}$ and $J_{\alpha}$. Indeed, we see in Fig. 15 that the error in $\Delta_{21}^{2}$ may reach to $\mathcal{O}\left(10\%\right)$ at $z\gtrsim 12$. At low redshifts ($z\lesssim 10$), the error due to this approximation goes to zero because at these redshift $\delta_{b}$ approaches more closely to $\delta_{c}$.

### v2

@@ Acknowledgements.
It is our pleasure to thank Julian B. Muñoz for reviewing the manuscript and making comments that improved the quality of this paper. We wish to thank the anonymous referee for providing a thorough and insightful report. We would also like to thank Andrei Mesinger and Tal Adi for useful discussions. We also acknowledge the efforts of the 21cmFAST and CLASS authors to produce state-of-the-art public 21-cm and CMB codes. JF is supported by the Zin fellowship awarded by the BGU Kreitmann School. SL is supported by an Azrieli International Postdoctoral Fellowship. EDK acknowledges joint support from the U.S.-Israel Bi-national Science Foundation (BSF, grant No. 2022743) and the U.S. National Science Foundation (NSF, grant No. 2307354), as well as support from the ISF-NSFC joint research program (grant No. 3156/23).
Appendix A Boltzmann equation for the spin temperature
@@
Combining Eqs. (A)-(41) thus yields
@@
Appendix B Scale-dependent matter evolution
As we discussed in Sec. II.5, the local SFRD can be computed with the EPS formalism, where the local linear matter over-density $\delta_{m}^{\mathrm{lin}}\left(\mathbf{x},z\right)$ is considered. This quantity can be computed in Fourier space via $\delta^{\mathrm{lin}}_{m}\left(\mathbf{k},z\right)\approx\mathcal{D}_{m}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where
$\sigma^{2}\left(M_{h},z\right)=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\mathcal{T}^{2}_{m}\left(k,z=0\right)\mathcal{D}_{m}^{2}\left(k,z\right)W^{2}\left(kR_{M_{h}}\right).$ (48)
Figure 10: Top panels: The ratio of $\sigma\left(M_{h},z\right)$ and the Sheth-Tormen HMF, when they are calculated with the SIGF and the SDGF. Bottom left panel: The ratio of the collapsed fraction of halos that host ACGs (with $M_{\mathrm{min}}$ given by Eq. V), when calculated with the SIGF and the SDGF. Bottom right panel: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the matter over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
Appendix C Scale-dependent Lagrangian perturbation theory
@@
C.1 First order - Zel’dovich approximation
@@
where we used the linear expression for $\delta_{i}$ (Eq. 13) as we seek first order solution to $S_{i}$. To see that this solution is consistent, we expand the Euler-Poisson equation, Eq. (56), to first order,
@@
which is indeed the equation of motion for $\delta_{i}$ in linear perturbation theory, above the Jeans scale (c.f. Eq. 11). Furthermore, the homogeneity assumption we made in Eq. (55) implies that $\delta_{i}$ and $S_{i}^{(1)}$ have the same zero initial conditions at $z=z_{0}$. Thus, the solution to $\mathbf{s}_{i}^{\left(1\right)}$ in Fourier space reads
@@ C.2 Second order - 2LPT
If variable separation was possible, then from Eq. (C.2) we could have obtained two different equations, one for the spatial part of $S_{i}^{(2)}$, and one for its temporal part. However, as we established in Eq. (61), variable separation cannot be performed for $S_{i}^{(1)}$, and thus $S_{i}^{(2)}$ cannot be separated to its spatial and temporal components either.
Note that given $S_{i}^{(1)}$, Eq. (C.2) does not give a differential equation for $S_{i}^{(2)}$, but rather for the Laplacian of $S_{i}^{(2)}$. In order to obtain an integro-differential equation for $S_{i}^{(2)}$, we take the Fourier transform of Eq. (C.2). While applying Fourier transform on the LHS gives a multiplication by $-k^{2}$, taking the Fourier transform of the RHS, being proportional to the square of second derivatives of $S_{i}^{(1)}$, would result in a convolution. We use the fact that in Fourier space $\ddot{S}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)+2H\dot{S}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)=4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(1\right)}\left(\mathbf{k},z\right)$ and apply the inverse Fourier transform to arrive at
@@
C.2.1 Scale independent growth
While in principle Eq. (C.2) can be solved numerically, this is computationally expensive, as for each cell $\mathbf{q}$ in the grid, double integration (that involves all the cells in the box in Fourier space) is required, for both species. To ease the computational cost, we will make the approximation that is done vastly in the literature, of approximating $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$, i.e., we replace the SDGF for each species with a universal SIGF. Then, from Eq. (61), we write
@@ C.2.1 Scale independent growth
where $\psi^{\left(1\right)}\left(\mathbf{q}\right)$ is the inverse Fourier transform of $\delta_{0}\left(\mathbf{k}\right)/k^{2}$ (or equivalently, in real space $\delta_{0}=-\nabla^{2}_{\mathbf{q}}\psi^{(1)}$). Because $\psi^{(1)}$ is derived from $\delta_{0}$, which is common for all species, under the SIGF approximation the solution to $S_{i}^{\left(1\right)}$ is a universal (i.e. independent of the species $i$) solution $S^{\left(1\right)}$. Plugging this universal solution back to Eq. (C.2) yields
@@
Because the RHS of Eq. (C.2.1) is universal for all species, and since $S_{i}^{(2)}$ has zero initial conditions at $z=z_{0}$ for every species $i$, we deduce that the SIGF approximation implies that $S_{i}^{(2)}\left(\mathbf{q},z\right)$ also has a universal solution $S^{(2)}\left(\mathbf{q},z\right)$. This property can be further inferred for any higher order in perturbation theory. Hence, the SIGF approximation corresponds to solving the perturbed equations of motions under the approximation $\delta_{b}\approx\delta_{c}\approx\delta_{m}$, namely we directly solve for the total matter density field.
After substituting the universal solution $S_{i}^{(2)}\left(\mathbf{q},z\right)=S^{(2)}\left(\mathbf{q},z\right)$, Eq. (C.2.1) can be written in the following form,
@@
C.2.2 Solution to $E\left(z\right)$
@@
Appendix D Temperature fluctuations in the dark ages
where $\kappa_{1-0}^{\mathrm{HH}}$ is the collision rate (in units of $\mathrm{cm^{3}/sec}$) of hydrogen atoms with themselves. Here, $\delta_{T}\left(k,z\right)$ and $\delta_{21,\mathrm{iso}}\left(k,z\right)$ are the Fourier transform of the fractional fluctuation in $T_{k}\left(\mathbf{x},z\right)$ and $T_{21}\left(\mathbf{x},z\right)$, respectively, and the latter does not contain fluctuations sourced from peculiar velocity (see Eq. 2). At the dark ages, the fluctuations $\delta_{T}$ are sourced by the perturbations in the baryon density field $\delta_{b}$, and thus it is useful to define the so called “adiabatic index”,
$k_{0}-\Delta k/2<|\mathbf{k}|<k_{0}+\Delta k/2,$ (89)